OpenStudy (anonymous):

6 people meet for a business lunch. Each person shakes hands once with each other person present. How many handshakes take place? A. 30 B. 21 C. 18 D. 15 E. 10

ganeshie8 (ganeshie8):

first person shakes hand wid remaining 5 ppl and leaves the lunch. = 5 hand shakes

ganeshie8 (ganeshie8):

second person shakes hand wid remaining 4 ppl and leaves the lunch. = 4 hand shakes

ganeshie8 (ganeshie8):

third person shakes hand wid remaining 3 ppl and leaves the lunch. = 3 hand shakes

ganeshie8 (ganeshie8):

.... total hand shakes = 5 + 4 + 3 + 2 + 1 = ?

OpenStudy (anonymous):

The equation is N(N-1)/2 N is 6. There are 6 people. (N-1) is 5. 6*5 = 30 30/2 = 15

OpenStudy (anonymous):

For problems like these. N number of people. Will shake hands with everybody else. (N-1) That is how we can derive. N(N-1). But Person 1 shaking hands with Person 2. And Person 2 shaking hands with Person 1 cannot be counted twice. Hence we divide by 2.

OpenStudy (anonymous):

It can also be written as nCr. N is the number of objects/entities. In this case 6. r is how many at a time. 2 people shake hands at a time. 6C2 = 15